OpenStudy (anonymous):

find the volume of the domain D: x^2+y^2≤2z ; 0≤z≤2

OpenStudy (chillout):

What kind of surface do you have?

OpenStudy (chillout):

\[\frac{x²}{a²}+\frac{y²}{b}= \frac{z}{c}\] What kind of surface is this?

OpenStudy (chillout):

I mean, over b²

OpenStudy (anonymous):

i dont know. i don't remember the formula. i've done this 3 years ago :)

OpenStudy (chillout):

It's a paraboloid... You can use double integrals so solve this. It intersects the plane z=2

OpenStudy (chillout):

Do you know how to do it?

OpenStudy (chillout):

I mean, the plane z = 4

OpenStudy (anonymous):

why a double and not a triple integral? \[\int\limits \int\limits \int\limits f(x,y,z) dxdydz\]

OpenStudy (chillout):

Just a moment, doing it on paper.

OpenStudy (anonymous):

http://clip2net.com/s/5efeON here is a print screen of a similar example. but i dont now what f i have here

OpenStudy (chillout):

Indeed it is supposed to be a triple integral. But it simplifies to a double one. Just one more moment...

OpenStudy (chillout):

Do you have the answer?

OpenStudy (anonymous):

no. it's the subject from last year exam.

OpenStudy (dan815):

want help?

OpenStudy (anonymous):

yes please :(

OpenStudy (dan815):

or u done?

OpenStudy (dan815):

ok

OpenStudy (dan815):

lets start by drawing

OpenStudy (dan815):

x^2+y^2≤2z z>x^2+y^2/2

OpenStudy (dan815):